PLEASE HELP MEH!

10. For the following question, consider a localized market for surface water with two competing parties. The city of West Chgiven by: -10 · Qs + 4000 if Q < 300 As an Inv. Demand Function: MBs= –5 · Qs + 2500 if Q> 300 As a Demand Function: G•)· MBs(b) (2 Points) What is the efficient total quantity of water in the market (Q5)? (c) (2 Points) What is the efficient margina(g) (2 Points) What quantity supplied is the cutoff below which water becomes scarce in this example?
Answer #1

A).ras am pudrops of Moolaitoineund Bro MB lano to 200 30 yoso Quempty

B).as market demand curve cuts supply curve at Q=350,so efficient quantity is 350.

C).as there is no external benefit or cost so,

Marginal social benefit=MARGINAL private benefit

=-5*350+2500=-1750+2500=750

D).mbi=mbi=750

E).qi=-1/10*750+400=-75+400=325

Qj=-1/10*750+100=-75+100=25

F). A good is scare when it available quantity or supply is less than its demand.

there is only 350 units supplied or available but entire demand is more than 350. So yes water is scare.

G).the cutoff is 500. As told in previous one ,that a good is scare if its maximum is lower then the entire demand.

So if water doesn't need to be scare ,it should be at least supply the maximum demand at any price.

The MAXIMUM demand is 500 ,so 500 is yhe cutoff of scarcity